# Shallow-KAN Based Solution of Moving Boundary PDEs

### v1

@@ Abstract
Kolmogorov–Arnold Networks (KANs) require significantly smaller architectures compared to multilayer perceptron (MLP)-based approaches, while retaining expressive power through spline-based activations. We propose a shallow KAN framework that directly approximates the temperature distribution $T(\textbf{x},t)$ and the moving interface $\Gamma(t)$, enforcing the governing PDEs, phase equilibrium, and Stefan condition through physics-informed residuals. To enhance accuracy, we employ interface-focused collocation resampling. Numerical experiments in one and two dimensions show that the framework achieves accurate reconstructions of both temperature fields and interface dynamics, highlighting the potential of KANs as a compact and efficient alternative for moving boundary PDEs. First, we validate the model with semi-infinite analytical solutions. Subsequently, the model is extended to 2D using a level-set based formulation for interface propagation, which is solved within the KAN framework. This work demonstrates that KANs are capable of solving complex moving boundary problems without the need for measurement data.
@@ 1 Introduction
Phase-change phenomena is ubiquitous and has applications in material processing [1, 2, 3], energy systems [4], biomedical [5, 6], geophysical [7]. Phase-change problems such as melting and solidification are commonly formulated as moving-boundary (Stefan-type) problems, where coupled transport PDEs determine the evolution of an interface separating phases [1, 8]. The interface (boundary dividing different phases) evolution is commonly modelled using a sharp [9, 10] or diffuse [11, 12] or hybrid-approaches [13]. For Stefan-type problems, stability and accuracy are tightly linked to how well the interface conditions and latent-heat balance are enforced; small temperature errors can induce large interface-location errors over time, motivating alternative space–time formulations beyond conventional discretisations[14, 15, 16].
@@
Physics-Informed Neural Networks (PINNs) [17] approximate the solution with a neural network trained by a loss composed of governing physics and are demonstrated to solve both forward and inverse problems [18, 19, 20]. Despite their appeal, PINNs face several challenges when extended to solve moving boundary problems. Firstly, they often have to incorporate measurement data to stabilize training, particularly in nonlinear heat transfer and two-phase Stefan problems [18]. Secondly, limited by spectral bias, standard PINNs struggle to resolve high-frequency variations [21, 22], which may affect the accuracy of gradients and sharp-boundary conditions on interfaces. Consequently, solving moving boundary PDEs generally requires deeper network architectures, which increases training cost and hampers the interpretability of the solution [23]. In a recent study  [24] investigated the use of PINN for phase transition problems, with adaptive weighting scheme and residual balancing and require substantial optimization effort and tuning. Another study by Mullins et al. [25] used MLP-type architecture for level-set equation to handle interfacial dynamics and demonstrated comparable performance similar to conventional techniques but may need careful tuning of weights and deep networks may demand very high samplig to extend it to coupled heat-transfer/stefan-type interface problems.
@@
In contrast to MLP based architectures, the recently proposed Kolmogorov-Arnold Networks (KANs) [26], grounded in the Kolmogorov-Arnold representation theorem [27, 28], have better interpretability and accuracy with shallower networks. They incorporate nonlinearity directly into learnable basis functions on the network edges rather than relying on fixed node activations, KANs with shallower networks matched comparable acurracy of standard MLP PINNs [26]. Recent studies have demonstrated that KANs effectively mitigate the spectral bias observed in standard MLP-based PINNs, enabling the efficient capture of fine-scale features and sharp gradients [29]. In this work, we propose a shallow KAN framework that approximates the temperature distribution and the moving interface by enforcing the governing PDEs, phase equilibrium (dirichlet condition of interface), and Stefan condition through physics-informed residuals. Interface capture is handled using Level Set method [30] adopted to KAN architecture.
@@ 2 Problem Statement & Methodology
The Stefan problem is a classical free-boundary problem modeling phase-change processes such as melting and solidification [31, 32]. Let $\Omega\subset\mathbb{R}^{3}$ denote the spatial domain, initially partitioned into a liquid region $\Omega_{\ell}(0)$ and a solid region $\Omega_{s}(0)$, separated by an interface $\Gamma(0)$. The objective is to determine the temperature fields $u_{\ell}(\mathbf{x},t)$ and $u_{s}(\mathbf{x},t)$ in $\Omega_{\ell}(t)$ and $\Omega_{s}(t)$, respectively, together with the evolution of the moving interface $\Gamma(t)$.
@@
and the interface motion is governed by the Stefan condition [32, 31]
@@
The unit normal vector pointing from solid to liquid is given by
$\mathbf{n}=\frac{\nabla\phi}{\|\nabla\phi\|},$ (9)
and the evolution of $\phi$ satisfies the level set advection equation,
$\phi_{t}+v_{n}\|\nabla\phi\|=0,$ (10)
where $v_{n}$ is the normal velocity of the interface.
@@
We now generalize the KAN-based solution framework for the $n$-dimensional two-phase Stefan problem. The model employs three KAN subnetworks to represent the unknown fields:
$(u_{s}(\mathbf{x},t),\;u_{\ell}(\mathbf{x},t),\;\phi(\mathbf{x},t)),$
where $u_{s}$ and $u_{\ell}$ are the temperature fields in the solid and liquid phases, and $\phi$ is the level set function defining the moving interface.
@@
where $H_{s}$ and $H_{\ell}$ denote smooth approximations of the characteristic functions of the solid and liquid regions, respectively. These masks are defined using a smoothed Heaviside function,
$H_{\ell}(\phi)$ $=\sigma(\beta\phi),$ (14)
$H_{s}(\phi)$ $=1-\sigma(\beta\phi),$ (15)
with $\sigma(z)=(1+e^{-z})^{-1}$ denoting the sigmoid function. Since $\phi(\mathbf{x},t)$ is the signed distance function from the interface, choosing a large value of $\beta$ (here $\beta=100$) yields a narrow transition zone, providing a smooth yet sharp approximation to the indicator functions of the liquid ($\phi>0$) and solid ($\phi<0$) domains.
@@
where the interface weighting function is chosen as a Gaussian centered at the zero level set,
$w_{\Gamma}(\phi)=\exp\!\left(-\frac{\phi^{2}}{\varepsilon_{\Gamma}^{2}}\right).$ (17)
The Gaussian weighting localizes the loss to a narrow band around the moving interface $\Gamma(t)=\{\mathbf{x}:\phi(\mathbf{x},t)=0\}$, providing a smooth and differentiable approximation to the sharp interface temperature continuity condition.
The level-set function $\phi(\mathbf{x},t)$ is defined as the signed normal distance from the solid–liquid interface, with $\phi>0$ in the liquid phase, $\phi<0$ in the solid phase, and $\phi=0$ on the interface $\Gamma(t)=\{\mathbf{x}:\phi(\mathbf{x},t)=0\}$. The interface is evolved by advecting $\phi$ according to
@@
Training points are generated as follows:
  • Collocation points: uniformly sampled throughout the spatio-temporal domain, with additional points concentrated near the interface $\Gamma(t)$ to enhance accuracy in regions with steep temperature gradients.
  • Boundary points: used to impose boundary conditions through $\mathcal{L}_{\text{BC}}$.
  • Initial points: used to satisfy $\mathcal{L}_{\text{IC}}$, defining the initial temperature distribution and initial interface location.
The training data generation method is summarized below.
@@
During optimization, gradient clipping is applied to avoid exploding gradients, and resampling of collocation points is performed iteratively to improve resolution near the moving interface. This adaptive resampling ensures that steep gradients at the solid–liquid boundary are accurately captured as training progresses. The overall training iteration for the n-dimensional Stefan problem is summarized below.
@@ 3 Results
In our experiments, a five-layer MLP with approximately 100 neurons per layer is required for each subnetwork $(u_{s},u_{\ell},s)$, resulting in roughly $1.2\times 10^{5}$ trainable parameters, whereas the corresponding KAN-based formulation uses only $640$ parameters.
@@
Figure 2: Evolution of the temperature field and error distribution for $n_{coll}=8000$ at $t=1.5$.
@@
500 $2.77\times 10^{-3}$ $3.40\times 10^{-5}$ 0.9990 $3.73\times 10^{-2}$ $1.68\times 10^{-3}$ 0.9992
1000 $2.43\times 10^{-3}$ $2.54\times 10^{-5}$ 0.9992 $2.90\times 10^{-2}$ $1.04\times 10^{-3}$ 0.9995
2000 $2.13\times 10^{-3}$ $2.08\times 10^{-5}$ 0.9994 $2.74\times 10^{-2}$ $9.49\times 10^{-4}$ 0.9996
4000 $1.79\times 10^{-3}$ $1.37\times 10^{-5}$ 0.9996 $2.31\times 10^{-2}$ $6.40\times 10^{-4}$ 0.9997
8000 $59\times 10^{-3} 1.1$$15\times 10^{-5}$ 0.9997 $1.93\times 10^{-2}$ $4.82\times 10^{-4}$ 0.9998
16000 $59\times 10^{-3} 1.1$$18\times 10^{-5}$ 0.9996 $1.99\times 10^{-2}$ $5.09\times 10^{-4}$ 0.9998
@@ 4 Conclusion
A key motivation for employing Kolmogorov–Arnold Networks (KANs) in the present work is their ability to represent nonlinear mappings with significantly fewer trainable parameters than conventional multilayer perceptrons (MLPs). KANs construct multivariate functions through compositions of adaptive univariate functions, which has been shown to improve parameter efficiency and reduce redundancy compared to standard fully connected architectures. While KANs have been used for solving PDEs, the present work specifically demonstrates their relevance and superiority in handling moving boundary problems, with interface jump conditions. This overcomes the existing limitation on employing measurement data. Further, we show that the shallow networks with two hidden layers and tens of learnable parameters are sufficient, even without using measurement data, in contrast to physics-informed MLP-based formulations involving nearly a million trainable parameters.

### v2

# A coupled Kolmogorov-Arnold Network and Level-Set framework for evolving interfaces

@@ Abstract
Kolmogorov–Arnold Networks (KANs) require significantly smaller architectures compared to multilayer perceptron (MLP)-based approaches, while retaining expressive power through spline-based activations. Moving interfaces are ubiquitous in physical systems, whose numerical solutions are quite complex. We propose a shallow KAN framework combined with a Level-set formulation that directly approximates the temperature distribution $T(\textbf{x},t)$ and the moving interface $\Gamma(t)$, enforcing the governing PDEs, phase equilibrium, and Stefan condition through physics-informed residuals. Numerical experiments in one and two dimensions show that the framework achieves accurate reconstructions of both temperature fields and interface dynamics, highlighting the potential of KANs as a compact and efficient alternative for moving boundary PDEs. First, we validate the model with semi-infinite analytical solutions. Subsequently, the model is extended to 2D using a level-set based formulation for interface propagation, which is solved within the KAN framework. This work demonstrates that KANs are capable of solving complex moving boundary problems without the need for measurement data.
@@ 1 Introduction
Phase-change problems such as melting and solidification are commonly formulated as moving-boundary (Stefan-type) problems, where coupled transport PDEs determine the evolution of an interface separating phases [1]. The interface (boundary dividing different phases) evolution is commonly modelled using a sharp [2, 3] or diffuse [4] or hybrid-approaches [5]. For Stefan-type problems, stability and accuracy are tightly linked to how well the interface conditions and latent-heat balance are enforced; small temperature errors can induce large interface-location errors over time, motivating alternative space–time formulations beyond conventional discretizations[6, 7].
@@
Physics-Informed Neural Networks (PINNs) [8] approximate the solution with a neural network trained by a loss composed of governing physics and are demonstrated to solve both forward and inverse problems [9, 10, 11]. Despite their appeal, PINNs face several challenges when extended to solve moving boundary problems. Firstly, they often have to incorporate measurement data to stabilize training, particularly in nonlinear heat transfer and two-phase Stefan problems [9]. Secondly, limited by spectral bias, standard PINNs struggle to resolve high-frequency variations [12, 13], which may affect the accuracy of gradients and sharp-boundary conditions on interfaces. Consequently, solving moving boundary PDEs generally requires deeper network architectures, which increases training cost and hampers the interpretability of the solution [14]. In a recent study  [15] investigated the use of PINN for phase transition problems, with adaptive weighting scheme and residual balancing and require substantial optimization effort and tuning.
Another study by Mullins et al. [16] used MLP-type architecture for level-set equation to handle interfacial dynamics and demonstrated comparable performance similar to conventional techniques but may need careful tuning of weights and deep networks may demand very high sampling to extend it to coupled Stefan-type interface problems.
@@
In contrast to MLP based architectures, the recently proposed Kolmogorov-Arnold Networks (KANs) [17], grounded in the Kolmogorov-Arnold representation theorem [18, 19], have better interpretability and accuracy with shallower networks. They incorporate nonlinearity directly into learnable basis functions on the network edges rather than relying on fixed node activations, KANs with shallower networks matched comparable acurracy of standard MLP PINNs [17]. Recent studies have demonstrated that KANs effectively mitigate the spectral bias observed in standard MLP-based PINNs, enabling the efficient capture of fine-scale features and sharp gradients [20]. In this work, we propose a shallow KAN framework that approximates the temperature distribution and the moving interface by enforcing the governing PDEs, phase equilibrium (dirichlet condition of interface), and Stefan condition through physics-informed residuals. Interface capture is handled using Level Set (LS) method [21] adopted to KAN architecture, termed KANLS methodology.
@@ 2 Problem Statement & Methodology
The Stefan problem is a classical free-boundary problem modeling phase-change processes such as melting and solidification. Let $\Omega\subset\mathbb{R}^{3}$ denote the spatial domain, initially partitioned into a liquid region $\Omega_{\ell}(0)$ and a solid region $\Omega_{s}(0)$, separated by an interface $\Gamma(0)$. The objective is to determine the temperature fields $u_{\ell}(\mathbf{x},t)$ and $u_{s}(\mathbf{x},t)$ in $\Omega_{\ell}(t)$ and $\Omega_{s}(t)$, respectively, together with the evolution of the moving interface $\Gamma(t)$.
@@
and the interface motion is governed by the Stefan condition
@@
[…]
@@
where $H_{s}$ and $H_{\ell}$ denote smooth approximations of the characteristic functions of the solid and liquid regions, respectively. These masks are defined using a scaled error function,
$H_{\ell}(\phi)$ $=\frac{1}{2}\left(1+\text{erf}\left(\frac{\phi}{\epsilon}\right)\right),$ (12a)
$H_{s}(\phi)$ $=\frac{1}{2}\left(1-\text{erf}\left(\frac{\phi}{\epsilon}\right)\right).$ (12b)
with $\text{erf}(z)=\frac{2}{\sqrt{\pi}}\int_{0}^{z}e^{-t^{2}}\,dt$ denoting the error function. Since $\phi(\mathbf{x},t)$ is the signed distance function from the interface, choosing a small value of $\epsilon$ (here $\epsilon=0.05$) yields a narrow transition zone, providing a smooth yet sharp approximation to the indicator functions of the liquid ($\phi>0$) and solid ($\phi<0$) domains.
@@
where the weighting function takes the form of the derivative of the scaled error function defined in Eq. (12).
$w_{\Gamma}(\phi)\propto\frac{dH}{d\phi}\implies w_{\Gamma}(\phi)=\exp\!\left(-\frac{\phi^{2}}{\epsilon^{2}}\right).$ (14)
The Gaussian weighting localizes the loss to a narrow band around the moving interface $\Gamma(t)=\{\mathbf{x}:\phi(\mathbf{x},t)=0\}$, providing a smooth and differentiable approximation to the sharp interface temperature condition. In order to capture the location of interface, given the interface velocity, the level-set function $\phi(\mathbf{x},t)$ is defined as the signed normal distance from the solid–liquid interface, with $\phi>0$ in the liquid phase, $\phi<0$ in the solid phase, and $\phi=0$ on the interface $\Gamma(t)=\{\mathbf{x}:\phi(\mathbf{x},t)=0\}$. The interface is evolved by advecting $\phi$ according to
@@
Training points are generated as follows: (i) Collocation points: uniformly sampled throughout the spatio-temporal domain, with additional points concentrated near the interface $\Gamma(t)$ to enhance accuracy in regions with steep temperature gradients, (ii) Boundary points: used to impose boundary conditions through $\mathcal{L}_{\text{BC}}$, (iii) Initial points: used to satisfy $\mathcal{L}_{\text{IC}}$, defining the initial temperature distribution and initial interface location.
The training data generation method is summarized below in Algorithm  1.
@@
During optimization, gradient clipping is applied to avoid exploding gradients, and resampling of collocation points is performed iteratively to improve resolution near the moving interface. This adaptive resampling ensures that steep gradients at the solid–liquid boundary are accurately captured as training progresses. The overall training iteration in the KANLS methodology for the n-dimensional Stefan problem is summarized below in Algorithm  2.
@@ 3 Results
For a problem of this nature, the architecture typically utilizes a five-layer MLP with approximately 100 neurons per layer for each of the subnetworks $(u_{s},u_{\ell},s)$, resulting in roughly $1.2\times 10^{5}$ trainable parameters [23], whereas the corresponding KAN-based formulation uses only $640$ parameters.
@@
(a) Exact solution
(b) KAN prediction
(c) Error map
Figure 2: Evolution of the temperature field and error distribution for $n_{\text{coll}}=8000$ at $t=1.5$.
@@
500 $3.74\times 10^{-3}$ $5.89\times 10^{-5}$ 0.9982 $4.18\times 10^{-2}$ $2.05\times 10^{-3}$ 0.9990
1000 $2.99\times 10^{-3}$ $3.68\times 10^{-5}$ 0.9988 $3.18\times 10^{-2}$ $1.22\times 10^{-3}$ 0.9994
2000 $2.76\times 10^{-3}$ $3.13\times 10^{-5}$ 0.9990 $2.99\times 10^{-2}$ $1.05\times 10^{-3}$ 0.9995
4000 $2.13\times 10^{-3}$ $1.71\times 10^{-5}$ 0.9995 $2.19\times 10^{-2}$ $5.46\times 10^{-4}$ 0.9997
8000 $1.47\times 10^{-3}$ $8.98\times 10^{-6}$ 0.9997 $1.54\times 10^{-2}$ $2.80\times 10^{-4}$ 0.9999
16000 $1.39\times 10^{-3}$ $8.43\times 10^{-6}$ 0.9997 $1.49\times 10^{-2}$ $2.79\times 10^{-4}$ 0.9999
@@ 4 Conclusion
A key motivation for employing Kolmogorov–Arnold Networks (KANs) in the present work is their ability to represent nonlinear mappings with significantly fewer trainable parameters than conventional multilayer perceptrons (MLPs). KANs construct multivariate functions through compositions of adaptive univariate functions, which has been shown to improve parameter efficiency and reduce redundancy compared to standard fully connected architectures. While KANs have been used for solving PDEs, the present work specifically demonstrates their relevance and superiority in handling moving boundary problems, with interface jump conditions. This overcomes the existing limitation on employing measurement data. Further, we show that the shallow networks with two hidden layers and tens of learnable parameters are sufficient, even without using measurement data, in contrast to physics-informed MLP-based formulations involving nearly a million trainable parameters. The KANLS methodology has a significant potential towards initiating further research in this field involving multi-phase physical systems.